Restart Night #9 UT 2018-09-10

OSA: D.G. Huerta (LBTO)

Observers/ISA: B. Rothberg, J. Hill, O.Kuhn (Tucson)

Software: S. Hooper/P. Kubanek (on-call)

ISp: J. Morris

Mtn Manager: R. Hansen

Details

UT 01:56 Some pesky clouds hanging around at the beginning to prevent us from opening.

UT 02:14 - Reconfiguring for MODS-Bino - will start with rotator zeropoint

UT 02:15 - MODS1 and 2 awake, running sievesnaps, LUCI software up and ready, LBC Blue trackers still not coming up

UT 02:19 - MODS 2 Red giving temp, LN2 warnings of 5C, RED IMCS failed - running blue only sievesnap

UT 02:24 - MODS2 cal lamp (VFLAT) did not turn on for sieve snap

UT 02:26 - retaking MODS2-Blue sievesnap - still nothing

UT 02:27 - manually turned on VFLAT (4) and waited 20sec - nothing

UT 02:30 - manually turned on VFLAT (8) - waited and took sievesnap - no flux

UT 02:33 - turned on QTH1 and took sievesnap - lots of light

UT 02:44 - updated Mask tables for both MODS - running massnap now - MODS2 Red giving dewar, CCD temp warning whenever scripts are run

Open

UT 03:15 - enclosure open, MODS in observing mode pointing and collimating check

MODS rotator center check

UT 03:30 - Stone O for rotator center check

UT 03:36 - Track Mode Stone O PA =0

mods1b.20180910.0005 mod1b.20180910.0005

mod2b.20180910.0009 mods2r.20180910.0004

UT 03:40 PA =180

mods1b 0006 mods1r 0006

mod2b 0010 mods2r 0005

UT 03:44 PA=90

mods1b 0007 mods1r 0007

mod2b 0011 mods2r 0006

UT 03:50 PA=270

mods1b 0008 mods1r 0008

mod2b 0012 mods2r 0007

Below added 1-Oct-2018 by OPK:

mods1b 5 & 6 -> 1532.6, 1542.6 and 7 & 8 -> 1533.8, 1547.2 avg value (1533,1545)

mods1r 5 & 6 -> 1528.6, 1547.5 and 7 & 8 -> 1530.6, 1551.9 avg value (1530,1550)

mods2b 9 & 10 -> 1508.3, 1504.5 and 1503.4,1511.2

MODS Rotator Zero Point Check

UT 03:55 - sending preset to stone O mask

UT 03:59 - had to alter MODS2 so it does acq in blue only (now a blue script in directory)

MODS-1 Alignment Result:

Computed Mask Alignment Offset:
dX = -0.494 arcsec
dY = 0.442 arcsec
dPA = -0.133 deg

MODS1 Offset Command:
offsetpointing -0.1333 -0.494 0.442 detxy rel

MODS-2 Alignment Result:

Computed Mask Alignment Offset:
dX = 4.040 arcsec
dY = -0.108 arcsec
dPA = 0.761 deg

MODS2 Offset Command:
offsetpointing 0.7609 4.040 -0.108 detxy rel

Alignment complete - will update Right Direct Rotator PCS Config

Spectrophotometric Standard Observation

UT 04:20 Feige 110

need to nudge MODS-1 6pixels left -0.74" (did this manually instead of modsAlign -r a 2nd time)

UT 04:30 Staring Feige 110 - cycling through all setups - Airmass = 1.774

UT 05:00 - blue PRISM MODS-1 0012--14

>> instconfig dual prism
STATUS: WARNING: **FAULT** - ROI Cannot change region-of-interest during acquiSTATUS: WARNING: **FAULT** - CCDBIN Data acquisition in progress. EXPSTATUS=INDONE: INSTCONFIG completed with 2 faults
>> imcslock
** ERROR: IMCSLOCK disallowed during exposures
** Abort, Retry, or Ignore? > R
** Retrying command 'imcslock'
appears to have remained at 8kX3k instead of changing to 4kX3k

Halting MODS-1 observations - everything after Dual Grating seems to have failed (race condition - raced ahead through script after IMCS lock to Prism)

mods1b 9-11 OK - ignore mods1b 12-14

mods1r 15-21 OK (1st red prism, #021 is ok)

MODS-1 did dual grating, red grating OK, choked at blue grating

all mods2blue data are OK (dual, blue, dual prism, blue prism)

UT 05:14 - Created temp script to pick up MODS-1 after Dual Grating

UT 05:15 - Blue IMCS MODS-1 having issues again

UT 05:17 - racing condition again, script continued even though IMCS lock was in progress, paused 1st exposure

Troubleshooting with Olga, tried sleep, wake, imcsTools on blue didn't show counts, then manually changed to blue and dual grating ran imcs lock manually that seemed to work

UT 05:47 - resending preset to take Dual Grating Feige 110 at lower airmass

UT 05:55 - Feige 110 airmass=1.368, mods1b 0015-0017, mods1r 0024-0026

UT 05:58 - MODS2 Blue IMCS lock failed and I hit another key after "R" and could not backspace so it aborted, created blue script for MODS-2 only mods2b 0033-0035

UT 06:14 - moving to Stone 0 GS=0 on axis for John Hill's collimation work

MODS AGw Tune-up

UT 05:24 In parallel to the standard star work, John updated the following parameters to make IDL, GCS and the actual MODS pupils all agree.

  • Changed MODS WFS_r_to_pupil_shift from 0.124 to 0.16 in MODS_L.cfg, MODS_R.cfg and IDL agw_define.pro for RDG. IDL for LDG already had the correct value of 0.16 which I've now propagated everywhere else.
  • Changed the IDL agw_define.pro RDG pupil hotspot from 400.6,225.6 to 402.6,222.6 to match what GCS MODS_R.cfg was already using. The GCS value was clearly better. With these two changes, all the MODS off-axis pupils are nicely centered on lenslet 7,7.
  • Change IDL LDG back to using the standard reconstructor from 20121213. It had previously been using 20130118 which had the leftmost column of lenslets removed.

With these three change IDL and GCS are in agreement about the corrections on both sides.

06:16 ACTIVE preset to STONE_O on-axis to check these AGW adjustments.

UT 06:19 - there is a transforminit.pro script in ~modseng/modsScripts/Support/

that configures MODS for transform collection. Need to make a version for Blue only MODS-2. Will run this first and then proceed with AGW Transform collection

putting pointing mask into place

IDL> iif_register,/TCS
IDL> iif_preset,Stone_O,mode=active,gs=0,pa=180,/left
IDL> iif_preset,Stone_O,mode=active,gs=0,pa=180,/right
IDL> iif_binoc_preset

Pointing Mask was not put into either MODS (got an error - and then manually verified that Slot 22 was empty - 3Kx3K image showed nothing)

Stone O for MODS Transform

Seeing has gotten bad and highly variable 1-1.4" - punting AGW transform for now

Moving to Vy 2-2 for 0.6" slit observations (bright target)

-60sec saturated a few lines in blue mods2b.20180910.0039, taking 15sec exposure mods2b.201809.0040 - worked for saturated lines

-60sec mods1r.20180910.0032 - halpha saturated, mods1b.20180910.0019 -few saturated

-taking 15sec exposures for mods1 - mods1r.20180910.0033, mods1b.20180910.0020 - halpha OK

-data looks OK - good enough for wavelength calibrations

Pointing Model Using Dave's PtStars IRAF program

manually configure MODS (imaging, filters, etc)

iif_register,/TCS
iif_preset,PT10m001,mode=active,gs=0,pa=0,/both
iif_binoc_preset
-on axis with GS=0

wait until collimate completes

home AGW, take exposure for sanity check (2sec)

/lbt/data/share/PointingLogs

20180910_SX.Log, 20180910_DX.Log

Current LST? (23.2):
Sorted in Azimuth:
AZ ALT RA J2000 DEC PtStarID MODS1b, MODS1r, MODS2b 20180910.00XX
======= ===================== ========
Vy2-2 19:24:22 09:53:56 - preceeding data had pointing info

8.19 50.08 00:19:22.047 71:35:43.13 PT10m001 022, 0035, 0042
9.53 73.65 23:28:13.610 48:46:01.30 PT10m101 (switched to 1Kx1K0) 0023, 0036, 0043
29.23 38.39 03:44:52.171 65:39:15.91 PT10m017 (landed where a cloud appeared) - no images taken but did collimate

UT 07:50 CLOSING


Taking MODS calibrations while we are closed

-Arc lamps Dual MODS-1, grpixflats Blue MODS-2 (mods1b 0024-0036 ; mods1r 0037-0039 ; mods2b 0044-0053)

timed out on Ne+Hg[Ar] lamps MODS-1 had to retry

-MODS-1 grpixflats Dual (mods1b 0027-0031; mods1r 0040--0049 )

-MODS-2 grlamps Blue (mods2b 0054-0056) - Ne didn't turn on for first exposure

-MODS 1 and MODS 2 8K biases (mods1b 0037-0041; mods1r 0045-0049; mods2b 0057-0061)

UT 08:31 - LBC finally came on! I ignored it so it finally worked

UT 08:34 - Taking 10 biases each LBC

UT 09:08 - taking slitflats Dual MODS-1, slitflats MODS-2 Blue (running only 0.6" - used tonight)

Need to verify that the new facility masks in MODS-2 are in the correct order - will take thru slits in Blue on both sides to compare

Use VFLAT 4 modsNX.20180910.YYYY - Names didn't get updated for first 3 on MODS-2 Blue, and first on MODS-1 Blue

MODS-1 Blue MODS-2 Blue Slit

0050 0069 0.3"

0048 0068 0.6"

0049 0070 0.8"

0051 0072 1.0"

0052 0071 1.2"

0053 0073 2.4"

UT 10:25 - calling summit

UT 10:40 - pointed and collimated near M76

UT 10:44 - doing test dithering and going through filters with MODS-1 and MODS-2 Blue imaging of M76 (first doing g,r MODS-1, u MODS-2)

MODS-1 Errors - instconfig failed - blue IMCS fail again?

[2018-09-10T10:49:57.141] MC1>OBS1 ERROR: IMCSLOCK RedIMCS =Fail BlueIMCS =Fail

UT 10:49 - stupidly resent preset on MODS-1 instead of continuing script - hopefully won't run into issues

Aborting both presets

mods2b.20180910.0074 u-sdss M76

MODS-1 had an issue where the name wasn't being set properly in setup

if you mistype in the MODS execBino windows that pop up, then you can't backspace and the whole thing crashes!!!!

Both Sides failed on

>> OBJNAME M76
** ERROR: Command 'OBJNAME M76' timed out after 60 seconds
Changed name to "Messier76"

Failed again on OBJNAME, changed it to "Target"

Tried script again, this time with mkMODSImg

FAILED AGAIN!!!!!!!!

UT 11:10 Presets hung??? Killed them

Trying imaging again -

Again, same error

* ERROR: Command 'OBJNAME M76' timed out after 60 seconds
*
Abort, Retry, or Ignore? >

Going to G191b2b spec std instead - this is ridiculous

Now Bino hangs on both for specstd

M1> ClearStars
** ERROR: Command 'ClearStars' timed out after 60 seconds
** Abort, Retry, or Ignore? >
UT 11:14 - Authorizing for MODS-1 only

This worked!!!!!

UT 11:19 - G191b2b

UT 11:25 - starting spectroscopic observations MODS-1 g191b2b Airmass=1.126

UT 11:37 - competed Spec Std G191B2b MODS-1

UT 11:43 - moving back to M76 imaging script that failed before - updated script so it cycles through all filters 4pt dither 30sec - seeing in guider and DIMM < 1"

mods1b.20180910.0057 -0068

mods1r.20180910.0062 - 0073

UT 12:12 - closing up, will take some MODS-1 Imaging Cals (flats and biases), putting MODS-2 to sleep, not planning on any cals

UT 12:20 - taking MODS-1 imaging flats

UT 12:25 - powering off LBCs

UT 01:00 - imaging flats done, 3K bias underway

UT 01:07 - MODS1r 3K bias 4/5 taking a long time to read out

UT 01:09 - putting MODS-1 to sleep

UT 01:10 - DONE!!!


TO DO

Need to update the PCSInstrument.conf in

/lbt/data/config/tcs/PCS
for AGW transform collection

-- %USERSIG{JohnHill - 2018-09-10}%
Topic revision: r11 - 02 Oct 2018, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback